OpenStudy (anonymous):

How do I prove that every powerful number can be written as the product of a perfect square and a perfect cube?

OpenStudy (anonymous):

Think about it, what do you know about a square?

OpenStudy (anonymous):

they are even

OpenStudy (anonymous):

there you go thats all it is

OpenStudy (kinggeorge):

Just curious, but what class is this for?

OpenStudy (asnaseer):

how does that prove it? I didn't know what "powerful numbers" were until I just looked them up. if I understand it correctly then a powerful number is a positive integer m such that for every prime number p dividing m, p^2 also divides m.

OpenStudy (anonymous):

Proofs class

OpenStudy (anonymous):

I didn't really understand the question

OpenStudy (asnaseer):

:/

OpenStudy (kinggeorge):

I was curious because I helped out on the same question yesterday. See http://openstudy.com/study#/updates/50552fcde4b02986d370aedd

OpenStudy (anonymous):

The first part makes sense but why are u subtracting 3 form ei when ei is odd?

OpenStudy (kinggeorge):

So that I get \[\large p_i^{e_i}=p_i^{f_i+3}=p_i^{f_i}\cdot p_i^3\]Note that since \(e_i\) is odd, \(e_i-3\) is even, so \(\displaystyle p_i^{e_i-3}=p_i^{f_i}\) is a perfect square.

OpenStudy (kinggeorge):

Additionally, \(p_i^3\) is a perfect cube.

OpenStudy (anonymous):

oh that makes perfect sense, thanks alot

OpenStudy (kinggeorge):

You're welcome.